OpenStudy (anonymous):

log(x)=-2 Solve for x?

OpenStudy (anonymous):

log(x)=-2

OpenStudy (ybarrap):

Assuming "log" is in base 10: $$ \large{ \log x=-2\\ 10^{\log x}=10^{-2}\\ x=\cfrac{1}{10^2}=\cfrac{1}{100}=0.01 } $$ Does this make sense?

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (ybarrap):

you're welcome